# DR2Seg: Decomposed Two-Stage Rollouts for Efficient Reasoning Segmentation in Multimodal Large Language Models

### v1

@@ 3.4 Theoretical Analysis
We analyze why self-rewarding with a two-stage rollout improves reinforcement learning for reasoning segmentation from both optimization and information-theoretic perspectives, compared to relying solely on answer-based rewards.
Optimization Analysis. For simplicity, we ignore the format reward and the non-repetition reward in Eq. 6. The objective of standard RL training based on answer rewards is given by:
@@
where $\mathcal{S}=(\mathcal{R},\mathcal{A})$ is the response with reasoning chains $\mathcal{R}$, and $\pi_{\theta}$ is the policy MLLM. Since $\mathcal{R}$ receives no direct supervision, the model tends to overthinking, producing excessively long reasoning that includes irrelevant content and sometimes drifts away from the intended reasoning objective. As a result, the uncertainty in training segmentation answers increases, leading to high-variance gradients.
We decompose the answer reward into two components: description reward and accuracy reward. The optimization objective is formulated as follows:
@@
where $\mathbf{R}_{\text{desc}}$ measures informativeness of the $\mathcal{D}$ and accuracy of the reasoning chains $\mathcal{R}$. This optimization considers both reasoning chains and spatial answers, providing direct signals for both multimodal reasoning and segmentation.
@@
This term measures how much information the reasoning chain provides for answer prediction beyond the given input image and question.
@@
Here, $\mathcal{D}$ serves as a compact semantic bottleneck that preserves task-relevant information while filtering redundant reasoning details. As a result, the conditional mutual information is reduced:
@@
making the second-stage reasoning more concise and stable.
Table 1: Performance comparison on the ReasonSeg benchmark. We additionally report the number of reasoning tokens to measure reasoning efficiency. * marks models trained on the train split of ReasonSeg. Bold and underlined values denote the best and second-best results, respectively. Our method shows notable superiority in both zero-shot and few-shot training settings.
@@ 4.1 Experimental Settings
Dataset. We first train the model using the 7K-sample setting of VisionReasoner (Liu et al., 2025b) for fair comparison, which is constructed from the LVIS (Gupta et al., 2019), RefCOCOg (Yu et al., 2016), gRefCOCO (Liu et al., 2023a), and LISA++ (Yang et al., 2023) datasets. We then follow the LISA (Lai et al., 2024) protocol and perform fine-tuning on the ReasonSeg train split, which contains only 239 samples with complex textual queries. For evaluation, we adopt the validation and test sets of ReasonSeg, and additionally report results on RefCOCO, RefCOCO+, and RefCOCOg to assess both common performance and generalization capability.
@@
Experimental Details. Following common practice (Liu et al., 2025b), we adopt Qwen2.5VL-7B (Bai et al., 2025) as the initial reasoning model and SAM2-Large (Ravi et al., ) as the segmentation model. Reinforcement learning is performed using the GRPO algorithm (Shao et al., 2024). Training is conducted with a total batch size of 16 with 8-sample rollout per training step. The initial learning rate is set to 1e-6, and the weight decay is 0.01. The predefined length anchor $N_{0}$ is set to 40, and the length penalty parameter $\gamma$ is set to 0.05. Following (Liu et al., 2025b), we train the model for one epoch on VisionReasoner-7K in the zero-shot setting, and for five epochs on the ReasonSeg training set in the few-shot setting. For all ablation studies, we train the model on the ReasonSeg training set and evaluate it on the ReasonSeg validation set. Notably, only the first-stage pass is used to ensure computational efficiency and a fair comparison during evaluation.
@@ 4.2 Main Results
Comparison methods. As shown in Table. 1 and Table. 2, the comparison methods are organized by columns: the first column lists non-MLLM methods (OVSeg (Liang et al., 2023), LAVT (Yang et al., 2022), ReLA (Liu et al., 2023a)); the second column includes SFT-based methods (LISA (Lai et al., 2024), CORES (Bao et al., 2024), PixelLM (Ren et al., 2024), Perception-GPT (Pi et al., 2024)); the third column presents representative RL-based methods (Seg-Zero (Liu et al., 2025a), SAM-R1 (Huang et al., 2025), PixelThink (Wang et al., 2025)); and the last column reports the recent state-of-the-art VisionReasoner (Liu et al., 2025b), which also serves as the baseline for our method, together with our DR2Seg.
@@
Referring Segmentation Results. We further report results on the referring expression segmentation (RES) task to evaluate the generalization ability of reasoning models in relatively simple scenarios, as shown in Tab. 2. Clearly, our DR2Seg method also demonstrates consistently strong performance, indicating that the decoupling strategy does not impair the reasoning model’s ability to handle simple scenarios. Furthermore, after fine-tuning on the ReasonSeg training set, DR2Seg* even improves upon DR2Seg, which is trained only on RES train datasets, further validating the generalization of our approach. Overall, DR2Seg* outperforms the baseline VisionReasoner*, achieving a 1.9% improvement on the refCOCOg dataset.
@@
Qualitative Results. Fig. 3 presents a qualitative comparison between VisionReasoner and DR2Seg. VisionReasoner exhibits an overthinking issue: despite correctly identifying the target object, excessive reasoning causes attention in during subsequent perception (e.g., the red-highlighted text misguides the MLLM to error regions). In contrast, DR2Seg reduces reasoning tokens and preserves focus on the target, achieving a better balance between efficiency and accuracy.
@@ 4.3 Diagnostic Experiments
Ablation on DR2Seg scheme. We ablate the two core self-rewards of DR2Seg in Tab. 3: the description self-reward $\mathbf{R}{\text{desc}}$ and the length-based self-reward $\mathbf{R}{\text{len}}$, both derived from our two-stage rollout strategy. By decoupling reasoning-based segmentation into multimodal reasoning and referring segmentation, the MLLM attains clearer task objectives and stronger focus on the target object. While $\mathbf{R}{\text{desc}}$ markedly improves accuracy, introducing $\mathbf{R}{\text{len}}$ further reduces reasoning tokens and even brings additional accuracy gains, validating the effectiveness of our two-stage design principles.
Effect of Two-stage Rollout.
To further analyze the effectiveness of the two-stage rollout, we examine the evolution of answer entropy, reasoning token count, and accuracy during training, as shown in Fig. 4. For answer entropy computation, we measure only the entropy of spatial answer tokens, excluding other reasoning tokens, to capture the MLLM’s uncertainty in target localization. Notably, the length-based reward is not applied in this experiment, allowing us to isolate the effect of the two-stage rollout structure. As training progresses, both answer entropy and reasoning token count consistently decrease, while accuracy steadily improves. These trends indicate that decoupling reasoning segmentation reduces overthinking, enables more confident target localization, and ultimately improves both accuracy and efficiency.
@@
Ablation on SAM3. We further evaluate our method on the latest segmentation model, SAM3 (Carion et al., 2025), which supports concept segmentation with brief phrases. Unlike SAM2, which directly supervises box and point predictions, we employ an external SAM3 API as a reward generator during training. Phrase descriptions produced by the MLLM are fed into SAM3 to obtain segmentation results, from which a IoU-based reward is computed (see supplementary material for details). Owing to the two-stage rollout strategy, our method naturally generates brief descriptions that align well with SAM3. As shown in Tab. 7, DR2Seg achieves significant improvements in both accuracy and efficiency, demonstrating its versatile.

### v2

@@ 3.4 Theoretical Analysis
In this section, we briefly analyze why this two-stage rollout improves RL-based reasoning segmentation from optimization and information-theoretic perspectives. The length-based reward is excluded to isolate the structural contribution. More details are provided in Sec. B of the Appendix.
Optimization Analysis. For clarity, we omit the format reward and the non-repetition reward in Eq. 6. Under standard reinforcement learning with answer-level supervision, the optimization objective is given by
@@
where $\mathcal{S}=(\mathcal{R},\mathcal{A})$ is the response with reasoning chains $\mathcal{R}$, and $\pi_{\theta}$ is the policy MLLM. Since $\mathcal{R}$ is not directly supervised, the model is prone to unconstrained exploration in the reasoning space, often resulting in excessively long reasoning chains. From an optimization perspective, this behavior increases the stochasticity of sampled trajectories, leading to high-variance gradients.
To alleviate this issue, we decompose the reward into two complementary components: a description reward and an answer reward. The resulting optimization objective is
@@
where $\mathbf{R}_{\text{desc}}$ provides intermediate supervision over descriptions $\mathcal{D}$, which are correlated with correct reasoning outcomes. By acting as an intermediate anchoring signal, $\mathbf{R}_{\text{desc}}$ improves credit assignment across the reasoning trajectory, thereby enabling more stable and efficient policy learning.
@@
This term measures how much information the reasoning chain provides for the answer beyond the given input image and question.
@@
Here, $\mathcal{D}$ serves as a compact information bottleneck that preserves necessary information. Assuming that $\mathcal{D}$ is a sufficient statistic extracted from $\mathcal{R}$ with respect to $\mathcal{A}$, the conditional mutual information is reduced:
@@
making the answer entropy of two-stage rollout reduced, as also supported by the experimental results in Fig. 4(a).
Table 1: Performance comparison on the ReasonSeg benchmark. We additionally report the number of reasoning tokens to measure reasoning efficiency. * marks models trained on the train split of ReasonSeg. Bold and underlined values denote the best and second-best results, respectively. Our method shows notable superiority in both zero-shot and few-shot settings.
@@ 4.1 Experimental Settings
Dataset. We first train the model using the 7K-sample setting of VisionReasoner (Liu et al., 2025b) for fair comparison, which is constructed from the LVIS (Gupta et al., 2019), RefCOCOg (Yu et al., 2016), gRefCOCO (Liu et al., 2023a), and LISA++ (Yang et al., 2023) datasets. We then follow the LISA (Lai et al., 2024) protocol and perform fine-tuning on the ReasonSeg train split, which contains only 239 samples with complex textual queries. For evaluation, we use the validation and test sets of ReasonSeg to evaluate performance in complex reasoning scenarios, and additionally report results on RefCOCO, RefCOCO+, and RefCOCOg to assess performance in referring segmentation.
@@
Experimental Details. Following common practice (Liu et al., 2025b), we adopt Qwen2.5VL-7B (Bai et al., 2025) as the reasoning model and SAM2-Large (Ravi et al., ) as the segmentation model. Reinforcement learning is performed using the GRPO algorithm (Shao et al., 2024). Training is conducted with a total batch size of 16 with 8-sample rollout per training step. The initial learning rate is set to 1e-6, and the weight decay is 0.01. The predefined length anchor $N_{0}$ is set to 45, and the length penalty parameter $\gamma$ is set to 0.05. Following (Liu et al., 2025b), we train the model for one epoch on VisionReasoner-7K in the zero-shot setting. We also train the model for five epochs on the ReasonSeg train set in the few-shot setting as (Lai et al., 2024). For all ablation studies, we train the model on the ReasonSeg train set and evaluate it on the ReasonSeg validation set. Notably, only the first-stage pass is used during evaluation to ensure computational efficiency and a fair comparison.
@@ 4.2 Main Results
Comparison Methods. As shown in Tab. 1 and Tab. 2, the comparison methods are organized by rows: the first row lists non-MLLM methods (OVSeg (Liang et al., 2023), LAVT (Yang et al., 2022), ReLA (Liu et al., 2023a)); the second row includes SFT-based methods (LISA (Lai et al., 2024), CORES (Bao et al., 2024), PixelLM (Ren et al., 2024), Perception-GPT (Pi et al., 2024)); the third row presents RL-based methods (Seg-Zero (Liu et al., 2025a), SAM-R1 (Huang et al., 2025), PixelThink (Wang et al., 2025)); and the last row reports the recent state-of-the-art VisionReasoner (Liu et al., 2025b), which also serves as the baseline for our method, together with our DR2Seg.
@@
Referring Segmentation Results. We further report results on the referring expression segmentation (RES) task to evaluate the generalization ability of reasoning models in relatively simple scenarios, as shown in Tab. 2. Clearly, our DR2Seg method also demonstrates consistently strong performance, indicating that the self-rewarding framework does not impair the model’s ability to handle simple scenarios. Furthermore, after fine-tuning on the ReasonSeg train set, DR2Seg* even improves upon DR2Seg, which is trained only on RES train datasets, further validating the generalization of our method. Overall, DR2Seg* outperforms the baseline VisionReasoner*, achieving a 1.9% improvement on the refCOCOg dataset.
@@
Qualitative Results. Fig. 3 presents a qualitative comparison between VisionReasoner and DR2Seg. VisionReasoner exhibits an overthinking issue: although it correctly identifies the target object, excessive reasoning leads to attention confusion in subsequent perception (e.g., the red-highlighted text misguides the MLLM toward erroneous regions). In contrast, DR2Seg reduces reasoning tokens while maintaining focus on the target, achieving a better balance between efficiency and accuracy. More qualitative results are provided in Sec. D of the Appendix.
@@ 4.3 Diagnostic Experiments
Ablation on DR2Seg scheme. We ablate the two core self-rewards of DR2Seg in Tab. 3: the description self-reward $\mathbf{R}{{}_{\text{desc}}}$ and the length-based self-reward $\mathbf{R}{{}_{\text{len}}}$, both derived from our two-stage rollout strategy. By decoupling reasoning segmentation into multimodal reasoning and referring segmentation, the MLLM attains clearer task objectives and stronger focus on the target object. $\mathbf{R}_{\text{desc}}$ leads to a notable improvement in accuracy, whereas the incorporation of $\mathbf{R}_{\text{len}}$ not only reduces the count of reasoning tokens but also yields an extra accuracy gain, thus validating the effectiveness of our two-stage design principle.
Effect of Two-stage Rollout. To further analyze the effectiveness of the two-stage rollout, we examine the evolution of answer entropy, reasoning token count, and accuracy during training, as shown in Fig. 4. For answer entropy computation, we measure only the entropy of segmentation answer tokens, excluding other reasoning tokens, to capture the MLLM’s uncertainty in localization. Notably, the length-based reward is not applied, allowing us to isolate the effect of the two-stage rollout structure. As training progresses, both answer entropy and reasoning token count consistently decrease, while accuracy steadily improves. These trends indicate that decoupling reasoning segmentation reduces overthinking, enables more confident target localization, and ultimately improves both accuracy and efficiency.
@@
Ablation on SAM3. We further evaluate our method on the recent segmentation model, SAM3 (Carion et al., 2025), which supports concept segmentation with brief phrases. Unlike SAM2 (Ravi et al., ), which directly supervises box and point predictions, we employ an external SAM3 API as a reward generator during training. Phrase descriptions produced by the MLLM are fed into SAM3 to obtain segmentation results, from which a IoU-based reward is computed (see Sec. C.3 of the Appendix). Owing to the two-stage rollout strategy, our method naturally generates brief descriptions that align well with SAM3. As shown in Tab. 7, DR2Seg achieves significant improvements in both accuracy and efficiency, demonstrating its versatility.